OpenStudy (anonymous):

What is 34% of 50? A. 7 B. 17 C. 34 D. 68

OpenStudy (anonymous):

I think it is 'B'

OpenStudy (usukidoll):

.34 x 50 =17 ya right

OpenStudy (anonymous):

Yay

OpenStudy (anonymous):

\[34\% \implies \frac{34}{100}\]\[\frac{34}{100} \cdot 50\]

OpenStudy (anonymous):

ok thank you